# Effect of spacetime geometry on neutrino oscillations

### v3

@@ I Introduction
Neutrinos in the Standard Model are massless and purely left-handed, described by two-component Weyl spinors. There is no right-handed neutrino in the Standard Model, as right-handed fermions are singlets of SU(2)${}_{L}$ and since neutrinos are colorless and electrically neutral, right-handed neutrinos are singlets of the full Standard Model gauge group and are thus must be decoupled from the rest of the theory. Absence of right-handed neutrinos means that there is no Dirac mass term, and neutrinos do not have a Majorana mass term because there is no Higgs triplet. On the other hand, the observation of disappearance (and in some cases, appearance) of flavors among atmospheric neutrinos [1], solar neutrinos [2], reactor neutrinos [3], accelerator neutrinos [4], and in many other experiments [5, 6, 7, 8, 9] point to neutrino flavor mixing and oscillation, which is explained by the existence of neutrino masses. This clearly indicates the existence of physics beyond the Standard Model.
@@
It goes without saying that before venturing into the unknown domain beyond the Standard Model, we should ensure that all known physics has been taken into account. Although the effects of Standard Model gauge fields on neutrino interactions are well studied, there is one sector of low energy physics which is rarely discussed in the context of neutrino oscillations, and that is gravitation – more precisely, spacetime geometry. Since gravity is a very weak force, it is generally thought that it will not contribute to neutrino processes. It was argued recently that there are ways to look at the contribution of spacetime geometry to neutrino oscillations which suggest that it may not be negligible in all situations [15]. The purpose of this paper is to make that suggestion more concrete, and possibly testable, by calculating such effects on neutrino conversion and survival probabilities.
@@ III Neutrino mixing and oscillation: Two neutrinos
We note here that other authors also have considered scenarios of neutrino oscillation with torsion [38, 39, 37, 40, 41, 42]. However, there are two important differences between these works and the model we have presented here: a) we do not consider a background torsion field, but only that which is generated by the presence of fermions, so it is a purely auxiliary field which can be integrated out; b) we consider a chiral coupling of the torsion to fermions, unlike in the existing literature.
@@ IV Neutrino mixing and oscillation: Three neutrinos
$i\frac{d}{dt}{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}$ $=\frac{\Delta\tilde{m}_{31}^{2}}{2E}\left[U^{*}\begin{pmatrix}0&0&0\\ 0&\alpha&0\\ 0&0&1\end{pmatrix}U^{T}+\begin{pmatrix}\tilde{A}&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}\right]{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}$ (81)
@@
$i\frac{d}{dx}{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}$ $=\frac{\Delta\tilde{m}_{31}^{2}}{2E}{\cal O}_{23}{\cal U}_{\delta}^{*}W\hat{M}W^{T}{\cal U}_{\delta}^{T}{\cal O}_{23}^{T}{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}$ (97)
@@ IV.1 Terrestrial conversion of $\nu_{\mu}$
[…]
@@
[…]

### v4

@@ I Introduction
Neutrinos in the Standard Model are massless and purely left-handed, described by two-component Weyl spinors. There is no right-handed neutrino in the Standard Model, as right-handed fermions are singlets of SU(2)L and since neutrinos are colorless and electrically neutral, right-handed neutrinos are singlets of the full Standard Model gauge group and are thus must be decoupled from the rest of the theory. Absence of right-handed neutrinos means that there is no Dirac mass term, and neutrinos do not have a Majorana mass term because there is no Higgs triplet. On the other hand, the observation of disappearance (and in some cases, appearance) of flavors among atmospheric neutrinos [1], solar neutrinos [2], reactor neutrinos [3], accelerator neutrinos [4], and in many other experiments [5, 6, 7, 8, 9] point to neutrino flavor mixing and oscillation, which is explained by the existence of neutrino masses. This clearly indicates the existence of physics beyond the Standard Model.
@@
It goes without saying that before venturing into the unknown domain beyond the Standard Model, we should ensure that all known physics has been taken into account. Although the effects of Standard Model gauge fields on neutrino interactions are well studied, there is one sector of low energy physics which is rarely discussed in the context of neutrino oscillations, and that is gravitation – more precisely, spacetime geometry. Since gravity is a very weak force, it is generally thought that it will not contribute to neutrino processes. However, it has long been suspected that the interplay of quantum mechanics with spacetime geometry may have an effect on neutrino oscillations. The gravitationally induced quantum mechanical phases discussed in [15, 16, 18, 19, 17] or the Berry phase induced by torsion, discussed in [20] are examples of this, as is the neutrino oscillation near black holes induced by a scalar conformally coupled to gravity [21]. Another class of examples come from considering the coupling of fermions to spacetime torsion [24, 25, 22, 23, 26]. Yet another possibility was suggested recently that dynamically generated spacetime torsion could provide another example where geometry has an observable effect on neutrino oscillations [27]. The purpose of this paper is to make that suggestion more concrete, and possibly testable, by calculating such effects on neutrino conversion and survival probabilities.
@@ III Neutrino mixing and oscillation: Two neutrinos
We have mentioned before that other authors have also considered scenarios of neutrino oscillation with torsion [22, 23, 24, 25]. However, there are three important differences between these works and the model we have presented here: a) we do not consider a background torsion field, but only that which is dynamically generated by the presence of fermions, so it is a purely auxiliary field which can be integrated out; b) we consider a chiral coupling of the torsion to fermions, unlike in the existing literature, and c) after integrating out the chiral torsion, we find a kind of MSW effect, but in the mass basis.
@@ IV Neutrino mixing and oscillation: Three neutrinos
$i\frac{d}{dt}{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}$ $=\frac{\Delta\tilde{m}_{31}^{2}}{2E}\left[U^{*}\begin{pmatrix}0&0&0\\ 0&\alpha&0\\ 0&0&1\end{pmatrix}U^{T}+\begin{pmatrix}\tilde{A}&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}\right]{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}$
@@
$i\frac{d}{dx}{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}$ $=\frac{\Delta\tilde{m}_{31}^{2}}{2E}{\cal O}_{23}{\cal U}_{\delta}^{*}W\hat{M}W^{T}{\cal U}_{\delta}^{T}{\cal O}_{23}^{T}{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}$
@@ IV.1 Terrestrial conversion of $\nu_{\mu}$
Let us consider oscillations of muon neutrinos into other flavors in the Earth’s crust which has a matter density $n$ of 2.7 gm/cm3 , for an energy range 0.5 GeV $<E<$ 10 GeV, which are appropriate for the DUNE experiment [59, 58, 61]. Also, only for the purpose of estimates, let us take all $\lambda_{f}$ and $\Delta\lambda_{ij}$ to be of the same order of magnitude, and write $\lambda^{2}$ to mean all quadratic combinations of the form $\lambda_{f}\Delta\lambda_{ij}\,.$ Then, for $\lambda^{2}\sim 0.1G_{F}$, we find that $\alpha$ is a small parameter of the same order of magnitude as $s_{13}$. Keeping terms up to the 2nd order in $\alpha$ and $s_{13}\,,$ we can write the amplitude of $\nu_{\mu}\to\nu_{\tau}$ conversion as [^4] For $\lambda^{2}\sim 1.0G_{F}\,,$ we cannot treat $\alpha$ as a small parameter, but we can numerically find the eigenvalues and the conversion and survival probabilities without using perturbation theory. We will do this below.
@@
As a demonstration, let us plot the conversion probabilities for neutrino oscillations in the DUNE experiment with and without the geometrical four-fermion interaction. While we do not know the size of the couplings $\lambda_{i}$ , we can guess that they are not very large compared to $\sqrt{G_{F}}$ , since otherwise there would be measurable differences with known results. Some guidance regarding the size of the couplings is provided by neutral current NSI bounds calculated in the literature, assuming that those take into account all possible types of four-fermion interactions. Let us plot the conversion probabilities when all the $\lambda_{f}$ are such that $O(\lambda_{f}^{2})=1G_{F}$ or less and compare with the case when they are set to zero. We have plotted the conversion probabilities $P(\nu_{\mu}\to\nu_{\tau})\,,P(\nu_{\mu}\to\nu_{e})\,,$ and the survival probability $P(\nu_{\mu}\to\nu_{\mu})\,,$ in Fig. 1, by numerically solving the Schrödinger equation. We have used the values [53] $\Delta m_{21}^{2}=7.39\times 10^{-23}~{}\text{GeV}^{2}\,,$ $\Delta m_{32}^{2}={{2.454\times 10^{-21}~{}\text{GeV}^{2}}}\,,\sin^{2}\theta_{23}=0.563\,,\sin^{2}\theta_{12}=0.31\,,\sin^{2}\theta_{13}=0.02237\,,\delta=221^{\circ}\,,$ and the average density of the Earth = 2.7 g/cm3 . We have also assumed that $\lambda$ for the electron, proton, and neutron are all of the same order of magnitude so that we can write $\tilde{n}\simeq\lambda n\,.$ The oscillation probabilities will then depend on quadratic combinations like $\lambda_{e}\Delta\lambda_{ij}$ , which we may write as $\lambda\Delta\lambda$ or as $\Delta\lambda^{2}$ for this purpose. As mentioned earlier, we have written $\lambda^{2}$ to represent the order of magnitude of all quantities quadratic in any $\lambda_{f}$ and $\Delta\lambda_{ij}$ In the plots, this magnitude is taken to be $\sim 0.1G_{F}\,$ and $\sim 1G_{F}\,,$ and we have plotted them along with the case $\lambda=0\,.$ The baseline $L$ is taken to be $1300~{}km$ and we have used normal mass ordering. We see from these plots that if the torsional coupling constants are as large as $\sqrt{G_{F}}\,,$ their effect on the conversion probabilities becomes quite large as the energy grows.
@@ Acknowledgements.
Data availability statement: This manuscript has no associated data or the data will not be deposited. [Authors’ comment: There are no data because this is a purely theoretical work.]